Problem 7 A NiCd battery has an amp-hour rating of 18.0 and consists of 13 cells, each with 1.2 V connected in series. What iHELP PLS!!!
Answer #1

Current I= 10 A. = = 9A zh ? - Now Max Power max = VxI = (13* 1-2 v) xq = 140-4 w Answer